OpenStudy (catearia):

A student runs 100 meters in 11 seconds. What is the speed of the student? speed = distance over time; 1 mile = 1609 meters

OpenStudy (wanell):

speed = distance / time speed = 100 / 11 speed = 9.09m/s to get rid of meters and get miles you want to multiply by 1 mile then divide by 1609 meters. to get hours, you times the amount of seconds in an hour (3600) and divide by 1 hour 9.09m/s × 1mile / 1609m × 3600s / 1 hour = 20.3 miles / hour

OpenStudy (wanell):

does that help